# Bidirectional Consistency Models

### v1

@@ Abstract
Diffusion models (DMs) are capable of generating remarkably high-quality samples by iteratively denoising a random vector, a process that corresponds to moving along the probability flow ordinary differential equation (PF ODE). Interestingly, DMs can also invert an input image to noise by moving backward along the PF ODE, a key operation for downstream tasks such as interpolation and image editing. However, the iterative nature of this process restricts its speed, hindering its broader application. Recently, Consistency Models (CMs) have emerged to address this challenge by approximating the integral of the PF ODE, thereby bypassing the need to iterate. Yet, the absence of an explicit ODE solver complicates the inversion process. To resolve this, we introduce the Bidirectional Consistency Model (BCM), which learns a single neural network that enables both forward and backward traversal along the PF ODE, efficiently unifying generation and inversion tasks within one framework. Notably, our proposed method enables one-step generation and inversion while also allowing the use of additional steps to enhance generation quality or reduce reconstruction error. Furthermore, by leveraging our model’s bidirectional consistency, we introduce a sampling strategy that can enhance FID while preserving the generated image content. We further showcase our model’s capabilities in several downstream tasks, such as interpolation and inpainting, and present demonstrations of potential applications, including blind restoration of compressed images and defending black-box adversarial attacks.
@@ 1 Introduction
Particularly, score-based diffusion models (DMs) [4, 5, 6, 7, 15] have stood out among their counterparts for generation [16]. Starting from random initial noise, DMs progressively remove noise through a process akin to a numerical ODE solver operating over the probability flow ordinary differential equation (PF ODE), as outlined by Song et al. [7]. However, while this iterative sampling procedure provides a natural trade-off between compute and sample quality [15], it slows down the generation process and hence limits broader applications.
This issue has been recently addressed by consistency models (CMs) [17, 18] which directly computes the integral from any given time step to the trajectory’s starting point. Building on top of CMs, Kim et al. [19] introduced the Consistency Trajectory Model (CTM), which estimates the integral between any two time steps along the PF ODE trajectory towards the denoising direction. Through these approaches, consistency model family enables image generation with only a single Number of Function Evaluation (NFE), while also offering a trade-off between speed and quality.
@@ 2 Background and Preliminary
Besides, in a follow-up work, Song and Dhariwal [18] suggested to use the Pseudo-Huber loss for $d$, along with other techniques that include setting $\mu=0$ (i.e., $\bm{\theta}^{-}\leftarrow\texttt{stopgrad}\left(\bm{\theta}\right)$), proposing a better scheduler function for $N$, adapting a better reweighting function $\lambda(t_{n})=\nicefrac{{1}}{{|t_{n}-t_{n+1}|}}$. Dubbed improved Consistency Training (iCT), these modifications significantly improve the performance. Therefore, unless otherwise stated, we inherit these techniques in our work.
@@ 3.3.2 Zigzag Sampling
According to Song and Dhariwal [18], the two-step zigzag sampler effectively reduces FID and is theoretically supported [25]. However, as Kim et al. [19] have highlighted, the process of injecting fresh noise introduces stochasticity, which can alter the content of the image with each iteration. We additionally reveal that if we fix the injected fresh noise in each iteration to enforce determinacy, the quality of the generated images significantly deteriorates.
Fortunately, our proposed BCM provides a direct solution to this dilemma, leveraging its capability to traverse both forward and backward along the PF ODE. Rather than manually reintroducing a large amount of fresh noise, we initially apply a small amount and let the network amplify it. Given the small scale of the initial noise, this method maintains a nearly deterministic sampling process, thereby leveraging the benefits of both fresh noise and determinacy. In a nutshell, for iteration $n$ ($n=N,N-1,\cdots,0$), we have
@@
We detail this scheme in Algorithm 4. Also, to verify its effectiveness, in Figure 2, we illustrate some examples to compare the generated images by 1) manually adding fresh noise, 2) manually adding fixed noise, and 3) our proposed sampling process, i.e., adding a small noise and amplifying it with the network. We can clearly see our method’s determinacy in the generated images.
@@ 4.1 Image Generation
            (a) 1-step
            (b) 2-step
            (c) 3-step
            (d) 4-step
Figure 3: Uncurated samples generated by BCM-deep. We can observe that our proposed sampling methods preserve the image context to a large degree.
@@ D.3 Coverage of Trajectory during Training
In Figure 11, we visualize the probability of selecting a $(t_{n},t_{n^{\prime}})$ pair during the BCT process, where $(t_{n},t_{n^{\prime}})$ is defined in Section 15. This offers insights into how well the entire trajectory is covered and trained. We can see that most of the probability mass is concentrated in small-time regions. This reveals some of our observations in the experiments:

### v2

@@ Abstract
Diffusion models (DMs) are capable of generating remarkably high-quality samples by iteratively denoising a random vector, a process that corresponds to moving along the probability flow ordinary differential equation (PF ODE). Interestingly, DMs can also invert an input image to noise by moving backward along the PF ODE, a key operation for downstream tasks such as interpolation and image editing. However, the iterative nature of this process restricts its speed, hindering its broader application. Recently, Consistency Models (CMs) have emerged to address this challenge by approximating the integral of the PF ODE, largely reducing the number of iterations. Yet, the absence of an explicit ODE solver complicates the inversion process. To resolve this, we introduce the Bidirectional Consistency Model (BCM), which learns a single neural network that enables both forward and backward traversal along the PF ODE, efficiently unifying generation and inversion tasks within one framework. Notably, our proposed method enables one-step generation and inversion while also allowing the use of additional steps to enhance generation quality or reduce reconstruction error. Furthermore, by leveraging our model’s bidirectional consistency, we introduce a sampling strategy that can enhance FID while preserving the generated image content. We further showcase our model’s capabilities in several downstream tasks, such as interpolation and inpainting, and present demonstrations of potential applications, including blind restoration of compressed images and defending black-box adversarial attacks.
@@ 1 Introduction
Particularly, score-based diffusion models (DMs) [4, 5, 6, 7, 15] have stood out among their counterparts for generation [16]. Starting from random initial noise, DMs progressively remove noise through a process akin to a numerical ODE solver operating over the probability flow ordinary differential equation (PF ODE), as outlined by Song et al. [7]. However, it typically takes hundreds of iterations to produce high-quality generation results. This slow generation process limits broader applications.
This issue has been recently addressed by consistency models (CMs) [17, 18] which directly computes the integral from any given time step to the PF ODE trajectory’s starting point. Building on top of CMs, Kim et al. [19] introduced the Consistency Trajectory Model (CTM), which estimates the integral between any two time steps along the trajectory towards the denoising direction. Through these approaches, the consistency model family enables image generation with only a single Number of Function Evaluation (NFE, i.e., number of network forward passes) while also offering a trade-off between speed and quality.
@@ 2 Background and Preliminary
Besides, in a follow-up work, Song and Dhariwal [18] suggested to use the Pseudo-Huber loss for $d$, along with other techniques that include setting $\mu=0$ (i.e., $\bm{\theta}^{-}\leftarrow\texttt{stopgrad}\left(\bm{\theta}\right)$), proposing a better scheduler function for $N$, adapting a better reweighting function $\lambda(t_{n})=\nicefrac{{1}}{{|t_{n}-t_{n+1}|}}$. Dubbed improved Consistency Training (iCT), these modifications significantly improve the performance. Therefore, unless otherwise stated, we inherit these improving techniques in our work.
@@ 3.3.2 Zigzag Sampling
According to Song and Dhariwal [18], the two-step zigzag sampler effectively reduces FID and is theoretically supported [25]. However, as Kim et al. [19] have highlighted, the process of injecting fresh noise can alter the content of the image after each iteration, which will be undesired, especially considering that our tasks will both include generation and inversion. One may immediately think that setting the injected noise to be the same as the initial random noise can fix this issue. However, we reveal that this will significantly damage the quality of the generated images.
Fortunately, our proposed BCM provides a direct solution to this dilemma, leveraging its capability to traverse both forward and backward along the PF ODE. Rather than manually reintroducing a large amount of fresh noise, we initially apply a small amount and let the network amplify it. Given the small scale of the initial noise, this method maintains a nearly deterministic sampling process, thereby leveraging the benefits of both fresh noise and determinacy. In a nutshell, for iteration $n$ ($n=N,N-1,\cdots,2$), we have
@@
where $\varepsilon_{n-1}$ is the scale of the small noises we add in $n$-th iteration. We detail this scheme in Algorithm 4. Also, to verify its effectiveness, in Figure 2, we illustrate some examples to compare the generated images by 1) manually adding fresh noise, 2) manually adding fixed noise, and 3) our proposed sampling process, i.e., adding a small noise and amplifying it with the network. We can clearly see our method maintains the generated content.
@@ 4.1 Image Generation
         (a) 1-step (FID 2.64)
         (b) 2-step (FID 2.36)
         (c) 3-step (FID 2.19)
         (d) 4-step (FID 2.07)
Figure 3: Uncurated samples generated by BCM-deep. We can observe that our proposed sampling methods can improve the FID while preserving the content to a large degree.
@@ D.3 Coverage of Trajectory during Training
In Figure 11, we visualize the probability of selecting a $(t_{n},t_{n^{\prime}})$ pair during the BCT process, where $(t_{n},t_{n^{\prime}})$ is defined in Equation 15. This offers insights into how well the entire trajectory is covered and trained. We can see that most of the probability mass is concentrated in small-time regions. This reveals some of our observations in the experiments: